how many miles of H2O will be produced from 2.9g of HCL reacting with Ca(OH)2


1
Expert's answer
2021-02-10T01:41:06-0500

Solution:

The balanced chemical equation:

Ca(OH)2 + 2HCl → CaCl2 + 2H2O

According to the equation above: n(HCl) = n(H2O)


Moles of HCl = Mass of HCl / Molar mass of HCl

The molar mass of HCl is 36.458 g mol-1.

Hence,

n(HCl) = 2.9 g / 36.458 g mol-1 = 0.07954 mol


n(H2O) = n(HCl) = 0.07954 mol

Moles of H2O = 0.07954 mol = 0.080 mol


Answer: About 0.080 moles of H2O will be produced.